# NaVIDA: Vision-Language Navigation with Inverse Dynamics Augmentation

### v1

@@ Abstract
Vision-and-Language Navigation (VLN) requires agents to interpret natural language instructions and act coherently in visually rich environments. However, most existing methods rely on reactive state–action mappings without explicitly modeling how actions causally transform subsequent visual observations. Lacking such vision–action causality, agents cannot anticipate the visual changes induced by its own actions, leading to unstable behaviors, weak generalization, and cumulative error along trajectory. To address these issues, we introduce NaVIDA (Navigation with Inverse Dynamics Augmentation), a unified VLN framework that couples policy learning with action-grounded visual dynamics and adaptive execution. NaVIDA augments training with chunk-based inverse-dynamics supervision to learn causal relationship between visual changes and corresponding actions. To structure this supervision and extend the effective planning range, NaVIDA employs hierarchical probabilistic action chunking (HPAC), which organizes trajectories into multi-step chunks and provides discriminative, longer-range visual-change cues. To further curb error accumulation and stabilize behavior at inference, an entropy-guided mechanism adaptively sets the execution horizon of action chunks. Extensive experiments show that NaVIDA achieves superior navigation performance compared to state-of-the-art methods with fewer parameters (3B vs. 8B). Real-world robot evaluations further validate the practical feasibility and effectiveness of our approach. Code and data will be available upon acceptance.
@@ 1 Introduction
Figure 1: Comparision of three VLN paradigms. (a) Vanilla VLN relies on state-action mapping, lacking the causality between vision and actions; (b) Forward Dynamics Modeling (FDM) adds an additional image generation task which predicts future observations but is compute-heavy and distractible; (c) NaVIDA uses inverse dynamics supervision to learn the causality between vision and actions with the action predicition task only, making the learning process more efficient and stable.
@@
Vision-and-Language Navigation (VLN) [3, 19, 32] is a challenging multimodal sequential decision-making task, requiring an agent to navigate effectively through a 3D visual environment based solely on natural language instructions. Accordingly, effective VLN requires not only precise instruction comprehension but also long-horizon temporal modeling and robust generalization to visually diverse, unseen scenes. Motivated by advances in video-based Multimodal Large Language Models (MLLMs) [4, 23, 48], recent studies [46, 45, 8, 39, 34, 49] have integrated these models into VLN to exploit their superior visual perception, semantic understanding, and cross-scene generalization.
Many existing works [46, 45, 8] typically formulate VLN as a short-sighted state-to-action reactive strategy. As shown in Fig. 1(a), they map current language instruction and past visual observations to a few subsequent actions. This simplification ignores the potential causal relationship between visual information changes and corresponding actions in navigation tasks: how current visual observations are transferred to the next visual observation through the influence of actions. The model cannot predict the visual consequences of its actions, thus failing to correct small biases that gradually accumulate into serious navigational instability, such as future viewpoint misalignment caused by slight premature turns. As shown in Fig. 1(b), recently some visual prediction-based works [34, 47, 16] combine forward dynamics modeling (FDM), attempting to use future observations as auxiliary signals for prediction, but such image generation tasks usually require additional prediction modules, such as diffusion-based world models [9, 6], which introduces additional computational burdens [31, 47] and causes the model to be confined to redundant details [35, 47].
To address these challenges, inspired by the inverse dynamic models (IDM), we propose NaVIDA, a VLN framework that integrates inverse-dynamics supervision (IDS) as an auxiliary signal to learn the causal relationship between visual changes and corresponding actions. This generates a language-independent visual motion concept, allowing the model to understand the physical consequences of its predicted actions without predicting dense image features, even if it cannot see the future frames. To provide IDS with richer and more discriminative learning signals, we introduce a hierarchical probabilistic action chunking (HPAC) mechanism that progressively merges low-level actions into multiple levels of action chunks. Specifically, we first group temporally adjacent atomic actions with consistent visual transition patterns into sub-chunks, and then further merge these sub-chunks into higher-level chunks, thus representing more complete and semantically coherent behavior. Such hierarchical mechanism allows the agent to predict over extended temporal horizons, leading to more stable navigation. Moreover, during inference, we propose an entropy-guided execution selection mechanism that adaptively adjusts the execution horizon of predicted action chunks according to their uncertainty, allowing for early replanning and timely trajectory correction. Extensive experiments demonstrate that NaVIDAsignificantly enhances the causal understanding of visual actions, achieving state-of-the-art (SOTA) performance on the VLN-CE [19] benchmarks with fewer parameters, thus establishing a simple and lightweight paradigm for VLN research.
@@
  • We propose NaVIDA, a novel VLN framework that explicitly integrates inverse-dynamics supervision to learn causal relationship between visual changes and corresponding actions, enabling stable navigation behaviors.
  • The NaVIDA adopts a HPAC mechanism to provide the inverse-dynamics supervision with richer and more discriminative learning signals, together with a dynamic entropy-guided execution horizon selection strategy to mitigate error accumulation during inference.
  • Comprehensive experiments on VLN-CE benchmarks demonstrate that NaVIDA attains SOTA performance with fewer parameters, confirming the efficiency and effectiveness of our design and highlighting the value of inverse dynamic modeling for VLN.
@@ 2 Related Work
Vision-and-Language Navigation. VLN tasks require an embodied agent to perceive, reason, and act in response to natural-language instructions. Early studies [15, 7, 22] focused on discrete navigation on connectivity graphs with panoramic RGB–D inputs, achieving strong in-domain results but relying on heavy sensors and pre-scanned viewpoints. Subsequent works [40, 19] have moved toward continuous control and low-level action prediction to better reflect real-world deployment, while several efforts [14, 36, 1, 11, 25, 30, 29] have investigated monocular RGB or RGB–D variants that relax reliance on panoramic sensing. Concurrently, the emergence of MLLMs has spurred monocular-only VLN systems that leverage vision–language backbones for stronger cross-scene transfer and lower hardware cost [46, 45, 8, 39]. Despite these advances, most VLN methods remain limited to reactive state-to-action mapping modeling, neglecting the potential causal relationship between changes in visual observation and corresponding actions. Lacking such causal grounding, agents struggle to anticipate visual consequences, which leads to unstable turning or stop behaviors and long-horizon error accumulation. In contrast, our NaVIDA explicitly encodes the action and induced visual change through inverse-dynamics supervision, enabling the policy to couple language grounding with short-horizon causal prediction and thereby stabilizing execution under monocular sensing.
@@
Dynamic Models in Navigation. Reactive state-action modeling directly maps historical visual observations and natural language instructions to the action space, lacking causal modeling between changes in visual information and actions. To mitigate this, early works [12, 21, 41] explored using additional pre-training tasks such as path-mask prediction, image prediction, or image completion. However, these were mostly limited to discrete environmental settings and often relied on extended contextual cues or language instructions. More recent methods [34, 47, 16] attempted to combine forward dynamics modeling and decompose the decision task, predicting future images while simultaneously predicting actions. Although effective in some settings, forward models typically require additional modules for high-dimensional prediction, increase compute and memory, and can bias learning toward redundant appearance details that are weakly aligned with the decision space. NaVIDA instead adopts inverse dynamics in a continuous environment: from two adjacent images it directly infers the action that caused the transition, sharing the output space with the reactive policy and avoiding extra generators or decoders. This design reduces optimization difficulty in mapping from image space to action space, improves data efficiency, and injects a causal prior that strengthens instruction grounding, yielding stable long-horizon navigation.
Figure 2: Architecture of NaVIDA. A multimodal language-model backbone outputs HPAC action blocks; training jointly optimizes navigation and IDM to strengthen the causal link between visual change and action, while inference uses entropy-guided execution horizon truncation to limit error accumulation.
@@ 3.1 Preliminary
Task Formulation. The Vision-and-Language Navigation task in continuous environment is defined as follows. At the beginning of each navigation episode, the agent receives a natural language instruction $\mathcal{I}$. Then, at each time step $t$, the agent receives an egocentric RGB observation $v_{t}$; we denote the observation history by $\mathcal{V}_{0:t}=\{v_{0},\ldots,v_{t}\}$. The discrete action space of the navigation agent is:
$\mathcal{A}=\{\textsc{Forward},\ \textsc{TurnLeft},\ \textsc{TurnRight},\ \textsc{Stop}\},$
and selecting an action $a_{t}\in\mathcal{A}$ triggers a transition to a new state and yields the next observation $v_{t+1}$. Episodes terminate when the agent issues Stop or reaches a maximum step budget $T_{\max}$. When the agent takes a stop action within 3 meters of the target position, the episode is considered successful.
Framework Overview. Motivated by the need to explicitly couple actions with the visual transitions they induce, we introduce NaVIDA on a standard VLM backbone consisting of a visual encoder, a vision–language projector, and an LLM head ( Fig. 2). On this backbone, our NaVIDA is organized around three mutually consistent components: (i) a variable-length action representation via Hierarchical Probabilistic Action Chunking (HPAC), (ii) inverse-dynamics supervision (IDS) injected at the data level during training, and (iii) an entropy-guided rule at inference to adaptively choose the executed prefix. The chunk grammar aligns IDS and VLN in the same action space; IDS strengthens vision-action causality without adding modules; entropy-guided execution stabilizes decoding by aligning horizon with model confidence.
@@ 3.2 Hierarchical Probabilistic Action Chunking
Instead of predicting a fixed small number of next actions, NaVIDA represents control as variable-length chunks to (a) amplify action-induced visual change for inverse-dynamic supervision and (b) allow farther-ahead decoding when confident, while keeping outputs consistent with the VLN action space. We use an HPAC mechanism to merge atomic actions in two stages. In the first probabilistic merging stage, up to $3$ temporally adjacent and visually consistent repetitions of the same atomic action are merged into a sub-chunk with probability $p$, which compacts micro-motions and reduces label noise. In the second, deterministic merging stage, up to $n$ consecutive sub-chunks are merged into a final chunk, yielding a sequence that covers between $n$ and $3n$ atomic actions. The resulting grammar preserves fine-grained short-range decisions while permitting farther-ahead decoding when confidence is high, and it amplifies action-induced appearance and viewpoint changes, thereby producing more discriminative supervision signals. The complete procedure is summarized in Algorithm 1. This chunk representation is used as the training target, providing a unified label space for IDS and navigation action prediction, and supporting chunk-level uncertainty estimation used by the entropy-guided execution rule.
Figure 3: IDM data construction pipeline. From Habitat trajectories, HPAC groups atomic actions into variable-length blocks and induces waypoints; start–end frames yield (current, goal) views, and the intervening actions are ground-truth labels.
@@ 3.3 Inverse-Dynamics Supervision
To reinforce the causal link between action and visual transition without modifying the architecture, we introduce inverse-dynamics supervision at the data level. Given an HPAC-induced pair of frames $(v_{t},v_{t+\ell})$, the model is required to predict the token sequence of the intervening variable-length action chunk $\mathbf{y}_{t}=(y_{t,1},\ldots,y_{t,n_{t}})$ using the same action grammar as VLN (i.e., action-type and numeric tokens). Supervision at the level of frame pairs and action chunks aggregates several micro-motions, leading to larger parallax and appearance changes and thus a more discriminative signal in settings with local aliasing (e.g., symmetry or transient occlusion). Following StreamVLN [39] and JanusVLN [43], we construct supervision data using Habitat [28] trajectories from R2R [3], RxR [20], ScaleVLN [38] and DAgger [27]. For each trajectory we use HPAC to define waypoints, and adjacent waypoints define non-overlapping subsequences, where the start frame is stored as the current view, the end frame as the goal view, and the intervening action chunk as the label ( Fig. 3). This yields approximately 2.42.4M IDS triplets and an additional $\sim\!2.8$M chunked VLN samples built under the same grammar.
@@
The combined objective $\mathcal{L}=\sum_{\tau\in D}\big(\mathcal{L}_{\mathrm{VLN}}^{\tau}+\mathcal{L}_{\mathrm{IDS}}^{\tau}\big)$ regularizes the representation to agree between actions that cause observed transitions and actions decoded under instructions, preserving the VLN action space and avoiding prediction of high-dimensional future appearance. With a unified chunk representation in place, inference can further exploit chunk-level uncertainty, which motivates the entropy-guided execution strategy described next.
@@
HPN+DN [17] ICCV2021 - $\checkmark$ $\checkmark$ $\checkmark$ 6.31 40.0 36.0 34.0
CMA [14] CVPR2022 - $\checkmark$ $\checkmark$ $\checkmark$ 6.20 52.0 41.0 36.0
@@
Sim2Sim [18] ECCV2022 - $\checkmark$ $\checkmark$ $\checkmark$ 6.07 52.0 43.0 36.0
GridMM [36] ICCV2023 - $\checkmark$ $\checkmark$ $\checkmark$ 5.11 61.0 49.0 41.0
ETPNav [2] TPAMI2024 - $\checkmark$ $\checkmark$ $\checkmark$ 4.71 65.0 57.0 49.0
ScaleVLN [38] ICCV2023 - $\checkmark$ $\checkmark$ $\checkmark$ 4.80 55.0 51.0
InstructNav [24] CoRL2024 - - - - - 6.89 31.0 24.0
LAW [26] EMNLP2021 - $\checkmark$ $\checkmark$ $\checkmark$ 6.83 44.0 35.0 31.0
CM2 [13] CVPR2022 - $\checkmark$ $\checkmark$ $\checkmark$ 7.02 41.5 34.3 27.6
ETPNav + FF [37] CoRL2024 - $\checkmark$ $\checkmark$ $\checkmark$ 5.95 55.8 44.9 30.4
Seq2Seq [19] ECCV2020 - $\checkmark$ $\checkmark$ 7.77 37.0 25.0 22.0
NavMorph [42] ICCV2025 - $\checkmark$ $\checkmark$ 5.75 56.9 47.9 33.2
NaVid [46] RSS2024 7B $\checkmark$ 5.47 49.1 37.4 35.9
NaVILA [8] RSS2025 8B $\checkmark$ 5.22 62.5 54.0 49.0
UniNaVid [45] RSS2025 7B $\checkmark$ 5.58 53.3 47.0 42.7
Aux-Think [33] NeurIPS2025 8B $\checkmark$ 6.08 60.0 54.8 46.9
StreamVLN [39] arXiv2507 7B $\checkmark$ 4.98 64.2 56.9 51.9
MonoDream [34] arXiv2508 2B $\checkmark$ 5.45 61.5 55.8 49.1
InternVLA-N1(S2) [10] - 7B $\checkmark$ 4.89 60.6 55.4 52.1
InternVLA-N1(S1+S2) [10] - 8B $\checkmark$ $\checkmark$ 4.83 63.3 58.2 54.0
NavFoM [44] arXiv2509 7B $\checkmark$ 5.01 64.9 56.2 51.2
JanusVLN [43] arXiv2509 8B $\checkmark$ 4.78 65.2 60.5 56.8
@@
3.4 Entropy-Guided Execution Horizon.
In multi-step decoding, executing the entire predicted sequence hampers responsiveness and accumulates errors, whereas executing too short a prefix may induce discontinuities and mode switching [32, 5]. To balance these effects, we leverage entropy as an uncertainty measure to assess the model’s confidence during action generation, and use it to adaptively truncate the execution horizon.
We calculate the action entropy of the model during inference. We ignore numeric tokens and focus solely on those representing action types, since once an action type is determined, the range of possible numerical tokens becomes highly constrained, and the choice of action type typically plays a more decisive role in navigation behavior at critical decision points. As is shown in Eq. 3, $N$ represents the number of action type tokens, $\mathcal{V}$ is the vocabulary and $p_{i}^{v}$ is the probability of the token $v$ at time $i$.
$H_{t}=-\sum_{i=1}^{N}\sum_{v\in\mathcal{V}}p_{i}^{v}\log p_{i}^{v}.$ (3)
We assume that the entropy of predicted actions should monotonically increase with the prediction horizon, reflecting the growing uncertainty of actions farther into the future. Conversely, a sudden entropy drop indicates inconsistency or overconfidence in the model’s long-horizon predictions. Since the final predicted action block typically corresponds to a distant and less reliable step, it is discarded. The sequence is truncated once the action entropy ceases to increase, ensuring that only the preceding action blocks with a consistent rising entropy trend are executed.
Formally, given a predicted action chunk list $\{a_{1},a_{2},\dots,a_{T}\}$ and their corresponding entropies $\{H_{1},H_{2},\dots,H_{T}\}$, the executed subset is defined as:
[…]
$\mathcal{A}_{\text{exec}}=\{\,a_{1},a_{2},\dots,a_{t^{*}}\,\}.$

### v2

@@ Abstract
Vision-and-Language Navigation (VLN) requires agents to interpret natural language instructions and act coherently in visually rich environments. However, most existing methods rely on reactive state-action mappings without explicitly action-grounded visual dynamics modeling. Lacking awareness of how actions transform subsequent visual observations, agents cannot plan actions rationally, leading to unstable behaviors, weak generalization, and cumulative error along trajectory. To address these issues, we introduce NaVIDA (Navigation with Inverse Dynamics Augmentation), a lightweight VLN framework that incorporates inverse dynamics supervision (IDS) as an explicit objective to embed action-grounded visual dynamics into policy learning. By jointly optimizing this visual dynamics with instruction-conditioned action prediction in a shared representation and action space, NaVIDA provides additional structured supervision that regularizes learning and leads to more stable and consistent navigation. To structure this supervision and extend the effective planning range, NaVIDA employs hierarchical probabilistic action chunking (HPAC), which organizes trajectories into multi-step chunks and provides discriminative, longer-range visual-change cues. Extensive experiments show that NaVIDA achieves superior navigation performance compared to state-of-the-art methods with fewer parameters (3B vs. 8B). Real-world robot evaluations further validate the practical feasibility and effectiveness of our approach. Code is available at https://github.com/waynechu1021/NAVIDA.
@@ 1 Introduction
Figure 1: Left: Comparision of three VLN paradigms. (a) Vanilla VLN relies on reactive state-action mapping, lacking the action-grounded visual dynamics; (b) Forward Dynamics Modeling (FDM) adds an additional image generation task which predicts future observations but is compute-heavy and distractible; (c) NaVIDA leverages inverse dynamics supervision to learn action-grounded visual dynamics through action prediction alone, resulting in more efficient and stable training. Right: Success rate versus model size comparison. NaVIDA (the blue one) achieves superior performance with smaller model size. The diameter of each circle is proportional to the model size (in billions of parameters).
@@
Vision-and-Language Navigation (VLN) [r2r, vlnce, vlnpe] is a challenging multimodal sequential decision-making task, requiring an agent to navigate effectively through a 3D visual environment based solely on natural language instructions. Accordingly, effective VLN requires not only precise instruction comprehension but also long-horizon temporal modeling and robust generalization to visually diverse, unseen scenes. Motivated by advances in video-based Multimodal Large Language Models (MLLMs) [qwen2_5, nvila, llava-video], recent studies [navid, uni-navid, navila, streamvln, monodream, zhang2025activevln] have integrated these models into VLN to exploit their superior visual perception, semantic understanding, and cross-scene generalization.
Many existing works [navid, uni-navid, navila] typically formulate VLN as a short-sighted state-to-action strategy. As shown in the left panel of Fig.˜1(a), this “Vanilla VLN” paradigm directly maps current language instructions and past visual observations to a limited set of subsequent actions. However, this reactive formulation lacks an explicit notion of action-grounded visual dynamics: how current visual observations are transferred to the next visual observation through the influence of actions. Without internalizing these visual dynamics, the agent merely learns shallow correlations between instructions and observations. Consequently, it cannot anticipate the visual consequences of the executed actions, rendering the navigation process highly susceptible to minor perceptual deviations, which inevitably accumulate over long horizons, leading to severe navigational instabilities, such as stopping too early or uncorrectable trajectory drift. To alleviate this issue, recent studies [monodream, zhang2025cross, huang2025vista] have explored Forward Dynamics Modeling (FDM, as illustrated in Fig.˜1b), which forces the agent to predict future visual observations. While FDM explicitly incorporates environmental dynamics into policy learning, it is inherently resource-intensive. Generating high-dimensional future frames often requires complex auxiliary modules (e.g., diffusion-based world models [chi2025diffusion, cai2025navdp]) and introduces significant computational overhead [wang2023dreamwalker, zhang2025cross]. More critically, FDM suffers from an inherent objective misalignment: reconstructing dense, pixel-level future frames forces the model to focus on redundant appearance details (e.g., lighting or textures) that are entirely irrelevant to navigation, rather than on optimizing the structured action space.
Recognizing that effective navigation does not require predicting every future pixel, we propose NaVIDA (left panel of Fig.˜1c), a VLN framework that incorporates inverse-dynamics supervision (IDS) as an auxiliary objective to introduce action-grounded visual dynamics into policy training. Rather than relying solely on instruction–action correlations, IDS trains the model to infer actions from visual transitions, encouraging consistency between predicted actions and the observable state changes they induce. This action-grounded supervision provides dense and language-agnostic training signals. By jointly aligning IDS with instruction-based action prediction within a shared feature and action space, NaVIDA facilitates the learning of visual motion representations that capture how actions drive environmental transitions—without requiring dense image prediction or access to future frames. To better structure this supervision and extend the effective planning horizon, we introduce a Hierarchical Probabilistic Action Chunking (HPAC) mechanism that progressively merges low-level actions into multi-level chunks. This hierarchical and stochastic merging strategy compacts short micro-motions into more informative training targets, providing richer and more discriminative signals for inverse dynamics learning. By introducing controlled randomness into chunk boundaries, HPAC avoids overly rigid compression, preserves segmentation diversity, and improves robustness to minor execution variations. As a result, the agent learns to reason over extended temporal spans while maintaining stable and consistent navigation behavior. Extensive experiments demonstrate the effectiveness of NaVIDA by achieving state-of-the-art (SOTA) performance on the VLN-CE [vlnce] benchmark with substantially fewer parameters (as visually summarized in Fig.˜1 Right), establishing a simple and lightweight paradigm for VLN research.
@@
  • We propose NaVIDA, a novel VLN framework that augments policy learning with action-grounded visual dynamics driven by inverse dynamics supervision (IDS). By aligning these two objectives in a shared feature and action space, NaVIDA achieves more stable and consistent navigation.
  • We introduce a Hierarchical Probabilistic Action Chunking (HPAC) mechanism to temporally structure the agent’s trajectory into variable-length semantic units. This design improves robustness to minor execution variations, and provides the inverse-dynamics objective with richer visual-transition cues, effectively extending the planning horizon.
  • Comprehensive experiments demonstrate that NaVIDA achieves SOTA performance with fewer parameters. Real-world tests confirm feasibility, robustness, and faster, more efficient execution, demonstrating the practical benefits of our joint action-understanding approach in physical environments.
@@ 2 Related Work
Vision-and-Language Navigation. VLN tasks require an embodied agent to perceive, reason, and act in response to natural-language instructions. Early studies [hong2021vln, chen2021history, liu2023bird] focused on discrete navigation on connectivity graphs with panoramic RGB-D inputs, achieving strong in-domain results but relying on heavy sensors and pre-scanned viewpoints. Subsequent works [wu2020towards, vlnce] have moved toward continuous control and low-level action prediction to better reflect real-world deployment, while several efforts [hong2022bridging, wang2023gridmm, an2022bevbert, nguyen2019vision, wang2022towards, schumann2024velma] have investigated monocular RGB or RGB-D variants that relax reliance on panoramic sensing. Concurrently, the emergence of MLLMs has spurred monocular-only VLN systems that leverage vision-language backbones for stronger cross-scene transfer and lower hardware cost [navid, uni-navid, navila, streamvln]. Despite these advances, most existing VLN methods remain limited to reactive state-to-action mappings, neglecting explicit modeling of how executed actions relate to subsequent visual transitions. Without such action understanding ability, agents struggle to anticipate short-horizon observation dynamics, often resulting in unstable turning, premature stopping, and compounding errors over long trajectories. In contrast, NaVIDA introduces inverse-dynamics supervision as an state-transition objective within policy learning. By aligning this with instruction-based action prediction in a shared feature and action space, our approach improves execution stability under monocular sensing.
@@
Dynamic Models in Navigation. Reactive state-action modeling directly maps historical visual observations and natural language instructions to the action space, without explicitly modeling the dependency between visual transitions and executed actions. To mitigate this, early works [mpmvln, vlnpredictfutureview, idmpretraininvln] explored using additional pre-training tasks such as path-mask prediction, image prediction, or image completion. However, these were mostly limited to discrete environmental settings and often relied on extended contextual cues or language instructions. More recent methods [monodream, zhang2025cross, huang2025vista] attempted to combine forward dynamics modeling and decompose the decision task, predicting future images while simultaneously predicting actions to enables modeling of state transitions. Although effective in some settings, forward models typically require additional modules for high-dimensional prediction, increase compute and memory, and can bias learning toward redundant appearance details that are weakly aligned with the decision space. NaVIDA instead adopts inverse dynamics in a continuous environment: from two adjacent images it directly infers the action that caused the transition, sharing the output space with the reactive policy and avoiding extra generators or decoders. This design reduces optimization difficulty in mapping from image space to action space, improves data efficiency, and injects an action-grounded transition prior constraints that strengthens instruction grounding, yielding stable long-horizon navigation.
Figure 2: Overview of NaVIDA. A multimodal language-model backbone generates HPAC action blocks. The HPAC mechanism stochastically merges adjacent actions into variable-length motion units, preserving chunk diversity and improves robustness to minor execution variations. Navigation action prediction and inverse dynamics supervision (IDS) are jointly optimized, enabling action-grounded visual dynamics modeling and strengthens action understanding within a shared representation and action space.
@@ 3.1 Preliminary
Task Formulation. The Vision-and-Language Navigation task in continuous environment is defined as follows. At the beginning of each navigation episode, the agent receives a natural language instruction $\mathcal{I}$. Then, at each time step $t$, the agent receives an egocentric RGB observation $v_{t}$; we denote the observation history by $\mathcal{V}_{0:t}=\{v_{0},\ldots,v_{t}\}$. The discrete action space of the navigation agent is: $\mathcal{A}=\{\textsc{Forward},\ \textsc{TurnLeft},\ \textsc{TurnRight},\ \textsc{Stop}\},$ and selecting an action $a_{t}\in\mathcal{A}$ triggers a transition to a new state and yields the next observation $v_{t+1}$. Episodes terminate when the agent issues Stop or reaches a maximum step budget $T_{\max}$. When the agent takes a stop action within 3 meters of the target position, the episode is considered successful.
Framework Overview. NaVIDA builds upon a standard Vision–Language Model (VLM) backbone, comprising a visual encoder, a vision–language projector, and an LLM head (Fig.˜2). NaVIDA jointly models instruction-based action prediction and action-grounded visual dynamics within a shared feature and action space. To provide structured and discriminative supervision signals, NaVIDA first introduces Hierarchical Probabilistic Action Chunking (HPAC), which organizes temporally adjacent actions into multi-level chunks and captures longer-range visual transition patterns. Built upon this hierarchical structuring, we incorporate inverse-dynamics supervision (IDS) as an auxiliary objective that trains the model to infer actions from visual transitions, strengthening action understanding and visual–action alignment beyond simple instruction–action correlations. Through the joint optimization of action prediction and IDS, NaVIDA achieves more stable and consistent navigation under monocular sensing.
@@ 3.2 Hierarchical Probabilistic Action Chunking
Instead of predicting a fixed small number of next actions, NaVIDA represents control as variable-length chunks to (a) amplify action-induced visual change for inverse-dynamic supervision and (b) allow farther-ahead decoding when confident, while keeping outputs consistent with the VLN action space. We use an HPAC mechanism to merge atomic actions in two stages. In the first probabilistic merging stage, up to $3$ temporally adjacent and visually consistent repetitions of the same atomic action are merged into a sub-chunk with probability $p$. This stochastic merging strategy compacts micro-motions and mitigates label noise, while avoiding the rigid and potentially over-aggressive compression induced by greedy schemes. By introducing controlled randomness into chunk boundaries, it preserves variability in action segmentation and improves robustness to minor execution fluctuations. In the second, deterministic merging stage, up to $n$ consecutive sub-chunks are merged into a final chunk, yielding a sequence that covers between $n$ and $3n$ atomic actions. The resulting grammar preserves fine-grained short-range decisions while permitting farther-ahead decoding when confidence is high, and it amplifies action-induced appearance and viewpoint changes, thereby producing more discriminative supervision signals. The complete procedure is summarized in Algorithm 1. This chunk representation is used as the training target, providing a unified label space for IDS and navigation action prediction.
@@ 3.3 Inverse-Dynamics Supervision
To enable action-grounded visual dynamics modeling without modifying the architecture, we introduce inverse-dynamics supervision (IDS) at the data level. Given an HPAC-induced pair of frames $(v_{t},v_{t+\ell})$, the model is required to predict the token sequence of the intervening variable-length action chunk $\mathbf{y}_{t}=(y_{t,1},\ldots,y_{t,n_{t}})$ using the same action grammar as VLN (i.e., action-type and numeric tokens). In this formulation, IDS serves as an explicit action-understanding objective, encouraging the model to interpret visual transitions in terms of structured action representations, while VLN remains an instruction-conditioned action-prediction task. By sharing the same feature and action space, the two objectives complement each other: IDS strengthens grounded action understanding from visual evidence, and VLN leverages this representation for language-guided decision making. Supervision at the level of frame pairs and action chunks aggregates multiple micro-motions, resulting in larger parallax and appearance variations and thus providing a more discriminative learning signal in visually ambiguous settings (e.g., symmetry or transient occlusion). Following StreamVLN [streamvln] and JanusVLN [janusvln], we construct supervision data using Habitat [habitat] trajectories from R2R [r2r], RxR [rxr], ScaleVLN [wang2023scaling], and DAgger [dagger]. For each trajectory, we apply HPAC to define waypoints, and adjacent waypoints determine non-overlapping subsequences, where the start frame is stored as the current view, the end frame as the goal view, and the intervening action chunk as the supervision label. This process yields approximately 2.42.4M IDS triplets and an additional $\sim\!2.8$M chunked VLN samples constructed under the same action grammar.
@@
The combined objective $\mathcal{L}=\sum_{\tau\in D}\big(\mathcal{L}_{\mathrm{VLN}}^{\tau}+\mathcal{L}_{\mathrm{IDS}}^{\tau}\big)$ regularizes the shared representation by encouraging consistency between action-grounded visual dynamics and instruction-conditioned action prediction. Actions inferred from observed transitions are aligned with actions decoded under language guidance, preserving the original VLN action space while avoiding the need to predict high-dimensional future visual appearance.
@@
HPN+DN [krantz2021waypoint] ICCV2021 - $\checkmark$ $\checkmark$ $\checkmark$ 6.31 40.0 36.0 34.0
CMA [hong2022bridging] CVPR2022 - $\checkmark$ $\checkmark$ $\checkmark$ 6.20 52.0 41.0 36.0
@@
Sim2Sim [krantz2022sim] ECCV2022 - $\checkmark$ $\checkmark$ $\checkmark$ 6.07 52.0 43.0 36.0
GridMM [wang2023gridmm] ICCV2023 - $\checkmark$ $\checkmark$ $\checkmark$ 5.11 61.0 49.0 41.0
ETPNav [an2024etpnav] TPAMI2024 - $\checkmark$ $\checkmark$ $\checkmark$ 4.71 65.0 57.0 49.0
ScaleVLN [wang2023scaling] ICCV2023 - $\checkmark$ $\checkmark$ $\checkmark$ 4.80 55.0 51.0
InstructNav [long2024instructnav] CoRL2024 - - - - - 6.89 31.0 24.0
LAW [law] EMNLP2021 - $\checkmark$ $\checkmark$ $\checkmark$ 6.83 44.0 35.0 31.0
CM2 [cm2] CVPR2022 - $\checkmark$ $\checkmark$ $\checkmark$ 7.02 41.5 34.3 27.6
ETPNav + FF [eptnavff] CoRL2024 - $\checkmark$ $\checkmark$ $\checkmark$ 5.95 55.8 44.9 30.4
Seq2Seq [vlnce] ECCV2020 - $\checkmark$ $\checkmark$ 7.77 37.0 25.0 22.0
NavMorph [navmorph] ICCV2025 - $\checkmark$ $\checkmark$ 5.75 56.9 47.9 33.2
NaVid [navid] RSS2024 7B $\checkmark$ 5.47 49.1 37.4 35.9
NaVILA [navila] RSS2025 8B $\checkmark$ 5.22 62.5 54.0 49.0
Uni-NaVid [uni-navid] RSS2025 7B $\checkmark$ 5.58 53.3 47.0 42.7
Aux-Think [aux-think] NeurIPS2025 8B $\checkmark$ 6.08 60.0 54.8 46.9
StreamVLN [streamvln] ICRA2026 7B $\checkmark$ 4.98 64.2 56.9 51.9
MonoDream [monodream] AAAI2026 2B $\checkmark$ 5.45 61.5 55.8 49.1
InternVLA-N1(S2) [internnav2025] - 7B $\checkmark$ 4.89 60.6 55.4 52.1
InternVLA-N1(S1+S2) [internnav2025] - 8B $\checkmark$ $\checkmark$ 4.83 63.3 58.2 54.0
NavFoM [navfom] ICLR2026 7B $\checkmark$ 5.01 64.9 56.2 51.2
JanusVLN [janusvln] ICLR2026 8B $\checkmark$ 4.78 65.2 60.5 56.8